OpenStudy (anonymous):

What is the distance between the points (3, 6) and (4, 2)?

OpenStudy (anonymous):

Do you know the distance formula?

OpenStudy (anonymous):

i don't know anything about math. its hard :/

OpenStudy (anonymous):

Distance between two pints is given as \[d=\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\] From the given points (3, 6) and (4, 2)we have \[x _{1}=3, x _{2}=4, y _{1}=6, x _{2}=2\] Plug these values in the above formula and u will be able to find the distance between the given points.

OpenStudy (anonymous):

Use the distance formula \[\sqrt{(x_{2}-x_1)^2+(y_2-y_1)^2}\]Insert (3, 6) and (4, 2). \[\sqrt{(4-3)^2+(2-6)^2}\]\[\sqrt{(1)^2+(-4)^2}\]\[\sqrt{(1+16)}\] \[\sqrt{(17)}\]That is the simplest form. \[\sqrt{(17)} \approx 4.1231056256\]

OpenStudy (anonymous):

Maths is very easy, need is to develop a small logical thinking.